OpenStudy (anonymous):

z= 10(cos45+iSin45) w= 5(cos15+iSin15) find (z)(w). How do I express this in polar form??

OpenStudy (anonymous):

is it a multiplication? Also, I believe the angle should be the same in one number.

OpenStudy (fibonaccichick666):

oh, I guess my response didn't go through, do you know euler's form?

OpenStudy (anonymous):

Yes Frouzen, it is multiplication. I went back and fixed the first sin value in the second equation.

OpenStudy (anonymous):

No, I do not know Euler's Form. What is it?

OpenStudy (fibonaccichick666):

This will help you immensely http://www.mathsisfun.com/algebra/eulers-formula.html

OpenStudy (fibonaccichick666):

The part where it is in the form \(r(cos\theta+isin\theta)\) is called the polar form of the complex number

OpenStudy (anonymous):

In the original formulas, where would the 10 and the 5 go in relation to Euler's form?

OpenStudy (fibonaccichick666):

Those are the radii of your circle

OpenStudy (fibonaccichick666):

scroll down towards the bottom of the link, it is well explained there

OpenStudy (fibonaccichick666):

so now, your mission is to convert each complex number to \(re^{i\theta}\) form do the multiplication, then take the result back into polar form